OpenStudy (alibaby):

WILL MEDAL Which of the following is a rational number? square root 1 square root 2, square root 3, square root 5 square root 1 square root 2 square root 03 square root 05

OpenStudy (alibaby):

@pooja195 @Will.H

OpenStudy (alibaby):

@callielovesyhuu

OpenStudy (alibaby):

@AloneS @Awsomesauce

OpenStudy (alibaby):

@HaileyD

jhonyy9 (jhonyy9):

Alibaby this your question was answered yesterday

jhonyy9 (jhonyy9):

do you know what is a rational number ?

jhonyy9 (jhonyy9):

a number is rational than can expressed in the form of a fraction for example 2 = 4/2 or 3 = 9/3 or 16 = 32/2 from these above wrote in your exercise what can be expressed sure easy in the form of a fraction ?

jhonyy9 (jhonyy9):

sqrt1 = 1 so in form of fraction how will be ?

OpenStudy (richhomiehelp):

@jhonyy9 what the answer for my question

OpenStudy (alibaby):

Ya ik but i got it wrong

OpenStudy (alibaby):

so im able to do it again

jhonyy9 (jhonyy9):

sqrt2 = 1,41 sqrt3 = 1,73

OpenStudy (alibaby):

i dont get it but i think the answer is c

jhonyy9 (jhonyy9):

why c. ?

jhonyy9 (jhonyy9):

sqrt3 = ? in fractional form ?

OpenStudy (alibaby):

nvm i think it's 1

OpenStudy (alibaby):

A*